# Avoiding Copyright Infringement via Large Language Model Unlearning

### v2

@@ Abstract
Pre-trained Large Language Models (LLMs) have demonstrated remarkable capabilities but also pose risks by learning and generating copyrighted material, leading to significant legal and ethical concerns. In real-world scenarios, model owners need to continuously address copyright infringement as new requests for content removal emerge at different time points. This leads to the need for sequential unlearning, where copyrighted content is removed sequentially as new requests arise. Despite its practical relevance, sequential unlearning in the context of copyright infringement has not been rigorously explored in existing literature. To address this gap, we propose Stable Sequential Unlearning (SSU), a novel framework designed to unlearn copyrighted content from LLMs over multiple time steps. Our approach works by identifying and removing specific weight updates in the model’s parameters that correspond to copyrighted content. We improve unlearning efficacy by introducing random labeling loss and ensuring the model retains its general-purpose knowledge by adjusting targeted parameters. Experimental results show that SSU achieves an effective trade-off between unlearning efficacy and general-purpose language abilities, outperforming existing baselines.  [^1] Code available at guangyaodou/SSU.
@@
Guangyao Dou1, Zheyuan Liu2, Qing Lyu1, Kaize Ding3, Eric Wong1 1University of Pennsylvania, 2University of Notre Dame, 3Northwestern University {gydou, lyuqing, exwong}@seas.upenn.edu, zliu29@nd.edu, kaize.ding@northwestern.edu
@@ 1 Introduction
In December 2023, the New York Times filed a lawsuit against OpenAI, accusing it of training its Large Language Models (LLMs) on copyrighted material without permission [^4] NYT Complaint, Dec 2023. This legal challenge highlights the growing concern over LLMs inadvertently incorporating copyrighted content from vast pre-training datasets, which are often composed of publicly available text  Min et al. (2023); Brittain (2023); Rahman and Santacana (2023). Despite the significant progress LLMs have made through learning from diverse text data  Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023), screening out copyrighted material remains an immense challenge Duarte et al. (2024). These issues raise broader questions about fair use of generative models.
@@
Previous works have investigated post-hoc copyright takedown methods – mitigating risks of generating copyrighted contents – using system prompt and decoding time intervention such as the MemFree Decoding Ippolito et al. (2022). Additionally, an alternative solution is machine unlearning Cao and Yang (2015), which removes unwanted knowledge after pre-training, reconfiguring the model as if it had never learned that data (Figure 3). Recent works proposed practical machine unlearning algorithms for LLMs Zhang et al. (2024b); Chen and Yang (2023); Jang et al. (2023); Zhao et al. (2024). However, few have addressed the challenge of sequentially unlearning copyrighted contents, leaving it unclear if existing methods are suitable as copyright takedown methods. An effective unlearning algorithm should be stable, meaning it should ensure unlearning efficacy while preserving non-targeted knowledge, books that are not subject to unlearning, and general-purpose language abilities.
@@
The core of many previous LLM unlearning methods have focused on Gradient Ascent (GA) and further combined it with an in-distribution retained dataset to preserve general-purpose language abilities, known as Gradient Difference Maini et al. (2024); Zhao et al. (2024); Liu et al. (2024d); Yao et al. (2024). However, Gradient Difference requires collection of a substantial amount of in-distribution retained data to maintain general-purpose abilities. Moreover, GA-based methods risk catastrophic collapse, where the model’s general-purpose language abilities degrade significantly Liu et al. (2024d). Zhang et al. (2024b) proposed Negative Preference Optimization (NPO), framing unlearning as preference optimization. However, NPO relies on a reference model, and if the reference model contains copyrighted information, unlearning efficacy is compromised.
@@
Our experiments on the Llama-3.1-8B-Instruct Dubey et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023) show that SSU achieves a superior trade-off between unlearning efficacy and general-purpose language abilities, avoiding the catastrophic collapse observed in GA-based methods. Moreover, SSU consistently outperforms NPO, which employs preference optimization frameworks. Additionally, fine-tuning with random loss and targeted model updates play distinct roles in facilitating the unlearning process within SSU. In contrast, copyright takedown methods that do not involve model weight updates, such as system prompts and MemFree Decode, fail to effectively mitigate the risks of copyright infringement.
@@ 2 Related Work
LLM Memorization and Copyright. LLMs are known to memorize and reproduce training data from the pre-training stage Carlini et al. (2021, 2022); Zhang et al. (2023); Nasr et al. (2023); Liu et al. (2024a). This behavior has prompted recent studies exploring the connection between verbatim memorization and copyright infringement Chu et al. (2024); Huang et al. (2024); Meeus et al. (2024); Karamolegkou et al. (2023), as well as the fair use of foundation models Henderson et al. (2023). Various methods have been proposed to mitigate these risks. For instance, Ippolito et al. (2022) introduced the MemFree decoding to reduce the likelihood of copyright infringement but failed to capture non-consecutive verbatim content. Additionally, Min et al. (2023) proposed SILO, a framework that utilizes a nonparametric datastore containing high-legal-risk data. However, the SILO framework does not address the risks associated with retrieval augmented generation Wei et al. (2024), nor does it adequately address the practical challenges of ensuring that the training data for parametric models is limited to permissive content.
Machine Unlearning. Machine Unlearning have been proposed for LLMs to remove harmful knowledge Liu et al. (2024d); Yao et al. (2023) and copyrighted content Jang et al. (2022); Eldan and Russinovich (2023); Liu et al. (2024b). Additionally, Maini et al. (2024) and Yao et al. (2024) explored the right to be forgotten and introduced benchmarks for evaluating the unlearning effectiveness of private data. However, none of these works have addressed the unlearning of copyrighted literary works in a sequential setting, nor the limitations of existing methods in this context.
Model Editing. A related but distinct process is model editing Meng et al. (2022); Hewitt et al. (2024); Gupta et al. (2024b). While unlearning aims to erase specific information from a model, model editing focuses on correcting errors without removing the underlying learned information Liu et al. (2024c). Recent studies have shown that model editing in sequential settings can lead to model collapse Gupta et al. (2024a); Yang et al. (2024), and potential solutions have been proposed to mitigate this issue Gupta and Anumanchipalli (2024).
@@ 3.2 Task Arithmetic
Unlearning via negating task vectors has recently gained attention Ilharco et al. (2022); Liu et al. (2024d) and has become an important baseline approach for many unlearning tasks. The rationale behind this approach is that by negating the gradient updates of the unwanted data, we can achieve a more localized unlearning algorithm to effectively erase $D_{f}$ from $\theta_{o}$.
@@ 4.1 Learning Stable Task Vectors
Random Labeling Loss. Inspired by previous works demonstrating that injecting noise during training improves learning outcomes Miyato et al. (2016); Srivastava et al. (2014); Neelakantan et al. (2015), we propose enhancing the effectiveness of unlearning by introducing data augmentation. Specifically, as shown in Figure 2 (b), we randomly mismatch the outputs of $D_{f}$ with the inputs of $D_{f}$. During the first stage of the task vector approach, we include the following loss:
@@
Weight Saliency. Moreover, to preserve general-purpose language abilities, we should mitigate the risk of catastrophic collapse during each time step of sequential unlearning. We can achieve this by steering the unlearning process towards specific parts of the model weights that are most relevant to the data to be forgotten. As shown in Figure 2 (c), we use a weight saliency map during the first stage of fine-tuning to further ensure localized unlearning by only adjusting specific weights that are most influenced by the data to be forgotten. The weight saliency map is defined as:
@@ 5.2 Evaluation Metrics
Figure 3: The averagd Rouge-1 and Rouge-L scores and benchmark scores for LLaMA3.1, omitting baseline methods that either consistently have low unlearning efficacy or easy to collapse (collapse details in Appendix C): (a) books to forget $D_{f}$ ($\downarrow$); (b) previously unlearned books $D_{prev}$ ($\downarrow$); (c) $D_{nor}$ ($\uparrow$). and (d) averaged normalized MMLU and MT-Bench scores ($\uparrow$). Lower Rouge scores on $D_{f}$ and $D_{prev}$ indicate better unlearning, while higher scores for $D_{nor}$ and benchmarks reflect better performance. Result with all methods is in Figure 6 in Appendix B.
@@ 5.4 Baseline Methods
System prompts. Following Wei et al. (2024), we use two system prompts to directly instruct the model to refrain from generating copyrighted content. The first, denoted as Prompting (a), adds instructions to not generate copyrighted contents to the default system prompt . The second, denoted as Prompting (dbrx), is used by the DBRX model from Databricks Team et al. (2024). This prompt specifies that the model was not trained on copyrighted books. It explicitly mentions that the model does not provide such content (More details can be found in Appendix A.1).
MemFree Decoding. The Memfree decoding checks whether the model’s next token would create an n-gram found in the Bloom filter Bloom (1970). If it would, we will choose the next most probable token and do the check again, until the generated n-gram will not be found by the filter.
Unlearning Methods. Unlearning methods such as Gradient Difference involves using books not in $D_{f}$ to maintain performance and a random mismatch loss to force random outputs for unlearned data. NPO treats $D_{f}$ as dispreferred data and formulates the unlearning problem within a preference optimization framework. More details, including hyperparameter choices, are in Appendix A.4.
@@ 6 Results
We present experimental results of Llama3.1 for different unlearning time steps in Figure 3 and 6. See the results for Mistral-7B in Appendix B, and the results with exact numbers in Appendix C.
@@ 6.1 Sequential Unlearning of Books
As shown in Figure 3(a) and 6(a), for Llama3.1, SSU consistently achieves one of the lowest average Rouge scores on $D_{f}$. Similarly, as shown in Figure 3(b) and 6(b), SSU proves to be one of the most effective across all time steps for books in $D_{prev}$, more effective in mitigating copyright infringement than NPO and Gradient Difference. The results are similar for Mistral-7B. As shown in Figures 7(a) and 7(b), Gradient Ascent and Gradient Difference achieve the lowest Rouge scores on $D_{f}$ at later time steps, effectively erasing copyrighted content through opposite gradient updates, ultimately leading to catastrophic collapse. TV also maintains effective unlearning, but also collapses at time step 5. In contrast, NPO’s average Rouge score on $D_{f}$ remains close to the vanilla model, indicating ineffectiveness in mitigating copyright infringement. On the other hand, SSU demonstrates less risks of copyright infringement.
@@
As shown in Figures 6(a) and 7(a), the Rouge scores for prompting and MemFree are largely indistinguishable from the vanilla model across many time steps and models. Although system prompts attempt to prevent generating copyrighted content, Llama3.1 often produces higher Rouge scores with prompting (a), and prompting (dbrx) is only marginally effective at certain time steps. We suspect this is due to the instruction-tuned Llama3.1 model’s inability to differentiate what constitutes copyrighted content. For the Mistral-7B model, both prompting methods are similarly ineffective in reducing infringement risks. Lastly, the MemFree decoding is always ineffective in unlearning copyrighted books for for Llama3.1 and Mistral-7B.
@@
Figure 4: Trade-off analysis between general-purpose language abilities and unlearning efficacy for Llama3.1 and Mistral-7B, considering only methods at time steps greater than 1. For improved visualization, we exclude Prompting (a) and GA, which consistently exhibit low unlearning efficacy or collapse during the process. We also exclude TV beyond time step 9 (Llama3.1) and time step 3 (Mistral-7B), as well as Gradient Difference beyond time step 4 (Mistral-7B), due to collapose at these stage (see Appendix C for collapse details). General-purpose abilities are calculated using normalized averages of MMLU and MT-Bench scores, while unlearning efficacy is represented by the average of Rouge-1 and Rouge-L scores on $D_{f}$ (targeted data) and $D_{prev}$ (previously unlearned data). Lower Rouge scores indicate better unlearning performance; hence, these values are negated for clarity. The ideal method balances both metrics and is positioned in the top-right corner. Full result with all methods is in Figure 8.
@@ 6.2 Non-Targeted Knowledge Retention
The results for Llama3.1 and Mistral-7B are notably different. For Mistral-7B, as shown in Figure 7(c), the average Rouge scores for GA, Gradient Difference, and TV demonstrate significant loss of retained knowledge at later time steps. While NPO is less effective at mitigating copyright infringement, it consistently retains more knowledge of non-targeted books than SSU. In contrast, the results for Llama3.1 fluctuate. As shown in Figures 3(c) and 6(c), Gradient Difference and SSU retain more knowledge of $D_{nrr}$ than the vanilla model. The unexpected re-emergence of these knowledge after unlearning is possibly due to knowledge redistribution during the unlearning process. Further research is needed to explore this phenomenon. Nevertheless, SSU demonstrates stability by preserving knowledge in $D_{nor}$.
@@
As shown in Figures 6(c) and 7(c), prompting and MemFree consistently retain non-targeted knowledge. Notably, for Llama3.1, both prompting (a) and MemFree retain more knowledge than the vanilla model. In contrast, for Mistral-7B, both prompting methods and MemFree exhibit slightly lower levels of knowledge retention compared to the vanilla model.

### v3

@@ Abstract
Pre-trained Large Language Models (LLMs) have demonstrated remarkable capabilities but also pose risks by learning and generating copyrighted material, leading to significant legal and ethical concerns. In real-world scenarios, model owners need to continuously address copyright infringement as new requests for content removal emerge at different time points. This leads to the need for sequential unlearning, where copyrighted content is removed sequentially as new requests arise. Despite its practical relevance, sequential unlearning in the context of copyright infringement has not been rigorously explored in existing literature. To address this gap, we propose Stable Sequential Unlearning (SSU), a novel framework designed to unlearn copyrighted content from LLMs over multiple time steps. Our approach works by identifying and removing specific weight updates in the model’s parameters that correspond to copyrighted content. We improve unlearning efficacy by introducing random labeling loss and ensuring the model retains its general-purpose knowledge by adjusting targeted parameters. Experimental results show that SSU achieves an effective trade-off between unlearning efficacy and general-purpose language abilities, outperforming existing baselines.  [^1] Code is available at guangyaodou/SSU.
@@
Guangyao Dou1  Zheyuan Liu2  Qing Lyu1  Kaize Ding3  Eric Wong1 1University of Pennsylvania  2University of Notre Dame  3Northwestern University gydou@seas.upenn.edu
@@ 1 Introduction
In December 2023, the New York Times filed a lawsuit against OpenAI, accusing it of training its Large Language Models (LLMs) on copyrighted material without permission [^4] NYT Complaint, Dec 2023. This legal challenge highlights the growing concern over LLMs incorporating copyrighted content from vast pre-training datasets, which are often composed of publicly available text  Brittain (2023); Rahman and Santacana (2023). Despite the significant progress LLMs have made through learning from diverse text data  Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023), screening out copyrighted material remains an immense challenge Duarte et al. (2024). These issues raise broader questions about fair use of generative models.
There are two times when copyright interacts with LLMs. The first is when LLMs learn from copyrighted materials, which is arguably fair use (but this has not been tested in court). The second is when LLMs generate outputs. If a generated output is substantially similar to copyrighted work it has trained on, then this is more likely to be a copyright infringement. If a court found an AI model developer to be in violation of copyright, then the court may require that the developer to remove that copyrighted work from the model. The cost of retraining from scratch leaving out one copyrighted work is exorbitantly high. Therefore, as an alternate remedy, a court may ask for a copyright takedown that does not require a full retraining of the model. This motivates research into unlearning and other copyright takedown methods.
@@
Previous works have investigated post-hoc copyright takedown methods – mitigating risks of generating copyrighted contents – using system prompt and decoding time intervention such as the MemFree Decoding Ippolito et al. (2022). Additionally, an alternative solution is machine unlearning Cao and Yang (2015), which removes unwanted knowledge after pre-training, reconfiguring the model as if it had never learned that data (Figure 3). Recent works proposed practical machine unlearning algorithms for LLMs Zhang et al. (2024); Chen and Yang (2023); Jang et al. (2023); Zhao et al. (2024). However, few have addressed the challenge of sequentially unlearning copyrighted content, where multiple unlearning requests must be processed over time without retraining from scratch. A key difficulty lies in controlling the changes in model weights during this process—existing methods often struggle to prevent unintended drift, leading to drastic degradation of general-purpose language abilities, leaving it unclear if they are suitable as copyright takedown methods. An effective unlearning algorithm should be stable, meaning it should ensure unlearning efficacy while preserving non-targeted knowledge, knowledge that are not subject to unlearning, and general-purpose abilities.
@@
The core of many previous LLM unlearning methods have focused on Gradient Ascent (GA) and further combined it with an in-distribution retained dataset to preserve general-purpose language abilities, known as Gradient Difference Maini et al. (2024); Zhao et al. (2024); Liu et al. (2024e); Yao et al. (2024). However, Gradient Difference requires collection of a substantial amount of in-distribution retained data to maintain general-purpose abilities. Moreover, GA-based methods risk catastrophic collapse, where the model’s general-purpose language abilities degrade significantly Liu et al. (2024e). Zhang et al. (2024) proposed Negative Preference Optimization (NPO), framing unlearning as preference optimization. However, NPO relies on a reference model, and if the reference model contains copyrighted information, unlearning efficacy is compromised.
@@
Our experiments on the Llama-3.1-8B-Instruct Dubey et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023) show that SSU achieves a superior trade-off between unlearning efficacy and general-purpose language abilities, avoiding the catastrophic collapse. Moreover, SSU consistently outperforms NPO, which employs preference optimization frameworks. Additionally, fine-tuning with random loss and targeted model updates play distinct roles in facilitating the unlearning process within SSU. In contrast, copyright takedown methods that do not involve model weight updates, such as system prompts and MemFree Decode, fail to mitigate the risks of copyright infringement.
@@ 2 Related Work
LLM Memorization and Copyright
LLMs are known to memorize and reproduce training data from the pre-training stage Carlini et al. (2021, 2022); Zhang et al. (2023); Nasr et al. (2023); Liu et al. (2024b). This behavior has prompted recent studies exploring the connection between verbatim memorization and copyright infringement Chu et al. (2024); Huang et al. (2024); Meeus et al. (2024); Karamolegkou et al. (2023), as well as the fair use of foundation models Henderson et al. (2023). Various methods have been proposed to mitigate these risks. For instance, Ippolito et al. (2022) introduced the MemFree decoding to reduce the likelihood of copyright infringement but failed to capture non-consecutive verbatim content. Additionally, Min et al. (2023) proposed SILO, a framework that utilizes a nonparametric datastore containing high-legal-risk data. However, the SILO framework does not address the risks associated with retrieval augmented generation Wei et al. (2024), nor does it adequately address the practical challenges of ensuring that the training data for parametric models is limited to permissive content.
Machine Unlearning
Machine Unlearning Cao and Yang (2015) have been proposed for LLMs to remove harmful knowledge Liu et al. (2024e); Yao et al. (2023), private information Dong et al. (2024), and copyrighted content Jang et al. (2022); Eldan and Russinovich (2023); Liu et al. (2024c). Additionally, Maini et al. (2024) and Yao et al. (2024) explored the right to be forgotten and introduced benchmarks for evaluating the unlearning effectiveness of private data. Li et al. (2024) further introduced a benchmark for forgetting hazardous knowledge in LLMs. However, none of these works have addressed the unlearning of copyrighted literary works in a sequential setting, nor the limitations of existing methods in this context.
Model Editing
A related but distinct process is model editing Meng et al. (2022); Hewitt et al. (2024); Gupta et al. (2024b). While unlearning aims to erase specific information from a model, model editing focuses on correcting errors without removing the underlying learned information Liu et al. (2024d). Recent studies have shown that model editing in sequential settings can lead to model collapse Gupta et al. (2024a); Yang et al. (2024a), and potential solutions have been proposed to mitigate this issue Gupta and Anumanchipalli (2024).
@@ 3.2 Task Arithmetic
Unlearning via negating task vectors has recently gained attention Ilharco et al. (2022); Liu et al. (2024e) and has become an important baseline approach for many unlearning tasks. The rationale behind this approach is that by negating the gradient updates of the unwanted data, we can achieve a more localized unlearning algorithm to effectively erase $D_{f}$ from $\theta_{o}$.
@@ 4.1 Learning Stable Task Vectors
Random Labeling Loss
Inspired by previous works demonstrating that injecting noise during training improves learning outcomes Miyato et al. (2016); Srivastava et al. (2014); Neelakantan et al. (2015), we propose enhancing the effectiveness of unlearning by introducing data augmentation. Specifically, as shown in Figure 2 (b), we randomly mismatch the outputs of $D_{f}$ with the inputs of $D_{f}$. During the first stage of the task vector approach, we include the following loss:
@@
Weight Saliency
While performing sequential unlearning, we implicitly leverage the knowledge of previous steps. Specifically, as new requests are introduced, the most recent unlearning step adjusts the model weights based on the parameters from earlier unlearning steps. One of the main challenges with existing methods is that they often fail to control the changes in model weights during this process, which can lead to instability and faster degradation of the model’s general-purpose capabilities.
Therefore, to preserve general-purpose language abilities, we should mitigate the risk of catastrophic collapse during each time step of sequential unlearning. We can achieve this by steering the unlearning process towards specific parts of the model weights that are most relevant to the data to be forgotten.
As shown in Figure 2 (c), we use a gradient-baased weight saliency map Fan et al. (2023) during the first stage of fine-tuning to further ensure localized unlearning by only adjusting specific weights that are most influenced by the data to be forgotten. The weight saliency map is defined as:
@@ 5.2 Evaluation Metrics
Figure 3: The averaged Rouge-1 and Rouge-L scores and benchmark scores for Llama3.1, omitting baseline methods that either consistently have low unlearning efficacy or easy to collapse (collapse details in Appendix D): (a) books to forget $D_{f}$ ($\downarrow$); (b) previously unlearned books $D_{prev}$ ($\downarrow$); (c) $D_{nor}$ ($\uparrow$). and (d) averaged normalized MMLU and MT-Bench scores ($\uparrow$). Lower Rouge scores on $D_{f}$ and $D_{prev}$ indicate better unlearning, while higher scores for $D_{nor}$ and benchmarks reflect better performance. Result with all methods is in Figure 6 in Appendix C.
@@ 5.4 Baseline Methods
System prompts
Following Wei et al. (2024), we use two system prompts to directly instruct the model to refrain from generating copyrighted content. The first, denoted as Prompting (a), adds instructions to not generate copyrighted contents to the default system prompt . The second, denoted as Prompting (dbrx), is used by the DBRX model from Databricks Team et al. (2024). This prompt specifies that the model was not trained on copyrighted books. It explicitly mentions that the model does not provide such content (More details can be found in Appendix A.1).
MemFree Decoding
The Memfree decoding checks whether the model’s next token would create an n-gram found in the Bloom filter Bloom (1970). If it would, we will choose the next most probable token and do the check again, until the generated n-gram will not be found by the filter.
Unlearning Methods
Unlearning methods such as Gradient Difference involves using books not in $D_{f}$ to maintain performance and a random mismatch loss to force random outputs for unlearned data. NPO treats $D_{f}$ as dispreferred data and formulates the unlearning problem within a preference optimization framework. More details, including hyperparameter choices, are in Appendix A.4.
@@ 6 Results
We present experimental results of Llama3.1 for different unlearning time steps in Figure 3 and 6. See the results for Mistral-7B in Appendix C, and the results with exact numbers in Appendix D.
@@ 6.1 Sequential Unlearning of Books
As shown in Figure 3(a) and 6(a), for Llama3.1, SSU consistently achieves one of the lowest average Rouge scores on $D_{f}$. Similarly, as shown in Figure 3(b) and 6(b), SSU proves to be one of the most effective across all time steps for books in $D_{prev}$, more effective in mitigating copyright infringement than NPO and Gradient Difference. The results are similar for Mistral-7B. As shown in Figures 9(a) and 9(b), Gradient Ascent and Gradient Difference achieve the lowest Rouge scores on $D_{f}$ at later time steps, effectively erasing copyrighted content through opposite gradient updates, ultimately leading to catastrophic collapse. TV also maintains effective unlearning, but also collapses at time step 5. In contrast, NPO’s average Rouge score on $D_{f}$ remains close to the vanilla model, indicating ineffectiveness in mitigating copyright infringement. On the other hand, SSU demonstrates less risks of copyright infringement.
@@
As shown in Figures 6(a) and 9(a), the Rouge scores for prompting and MemFree are largely indistinguishable from the vanilla model across many time steps and models. Although system prompts attempt to prevent generating copyrighted content, Llama3.1 often produces higher Rouge scores with prompting (a), and prompting (dbrx) is only marginally effective at certain time steps. We suspect this is due to the instruction-tuned Llama3.1 model’s inability to differentiate what constitutes copyrighted content. For the Mistral-7B model, both prompting methods are similarly ineffective in reducing infringement risks. Lastly, the MemFree decoding is always ineffective in unlearning copyrighted books for for Llama3.1 and Mistral-7B.
@@
Figure 4: Trade-off analysis between general-purpose language abilities and unlearning efficacy for Llama3.1 and Mistral-7B, considering only methods at time steps greater than 1. For improved visualization, we exclude Prompting (a) and GA, which consistently exhibit low unlearning efficacy or collapse during the process. We also exclude TV beyond time step 9 (Llama3.1) and time step 3 (Mistral-7B), as well as Gradient Difference beyond time step 4 (Mistral-7B), due to collapose at these stage (see Appendix D for collapse details). General-purpose abilities are calculated using normalized averages of MMLU and MT-Bench scores, while unlearning efficacy is represented by the average of Rouge-1 and Rouge-L scores on $D_{f}$ (targeted data) and $D_{prev}$ (previously unlearned data). Lower Rouge scores indicate better unlearning performance; hence, these values are negated for clarity. The ideal method balances both metrics and is positioned in the top-right corner. Full result with all methods is in Figure 10.
@@ 6.2 Non-Targeted Knowledge Retention
The results for Llama3.1 and Mistral-7B are notably different. For Mistral-7B, as shown in Figure 9(c), the average Rouge scores for GA, Gradient Difference, and TV demonstrate significant loss of retained knowledge at later time steps. While NPO is less effective at mitigating copyright infringement, it consistently retains more knowledge of non-targeted books than SSU. In contrast, the results for Llama3.1 fluctuate. As shown in Figures 3(c) and 6(c), Gradient Difference and SSU retain more knowledge of $D_{nor}$ than the vanilla model. The unexpected re-emergence of these knowledge after unlearning is possibly due to knowledge redistribution during the unlearning process. Yang et al. (2024b) also examined this anticipatory recovery behavior where LLMs recover knowledge from the forgetting on documents before encountering them again. Further research is needed to explore this phenomenon. Nevertheless, SSU demonstrates stability by preserving knowledge in $D_{nor}$.
@@
As shown in Figures 6(c) and 9(c), prompting and MemFree consistently retain non-targeted knowledge. Notably, for Llama3.1, both prompting (a) and MemFree retain more knowledge than the vanilla model. In contrast, for Mistral-7B, both prompting methods and MemFree exhibit slightly lower levels of knowledge retention compared to the vanilla model.